OpenStudy (anonymous):

8z-8-2z+2=0 plz help!!! i dont get itt!! i need the answer

OpenStudy (anonymous):

6z-6=0 z-1=0 z=1

OpenStudy (anonymous):

@mnarin The trick is getting anything with a letter after it to be alone...they don't like regular numbers :-) so get all of the regular numbers on one side of the equation and all of the numbers with letters (variables) on the other.

OpenStudy (anonymous):

thank u thats helpful! ugh.. i have a whole math quiz to do tonight by midnight and i dont know how to do polynomials!!

OpenStudy (anonymous):

To get rid of the -8 on the left, add a positive 8 to the right. Alternately, get rid of the positive 2 on the left by adding a negative 2 to the right hand side.

OpenStudy (anonymous):

You end up with 8z-2z on the lft...reduce that down to 6z as sheg wrote

OpenStudy (anonymous):

so the answer would then become 8z-2z 6z?

OpenStudy (anonymous):

add 8z and -2z 8z-2z=6z then add -8 and 2 -8+2=-6 now add these two 6z-6=0 subtract 6 from both sides 6z=6 divide both sides by 6 z=1

OpenStudy (anonymous):

you are AMAZINGG!!! :) Im gunna be posting more math questions cause ive got 20 i will need help for.

OpenStudy (anonymous):

no problem. i'll answer them all if you'd like

OpenStudy (anonymous):

lol your great! thank you so so so much! its for my midterm and i am so glad u can help me out. ill start the test and we can get started. when im done i hope this site can copy n paste to word so i can learn from it and study up on this.

OpenStudy (anonymous):

ok first one..

OpenStudy (anonymous):

(4u exponent 4 over 8v exponent 3) exponent 3 =

OpenStudy (anonymous):

sorry i have to type it out cause i dunno how to do it the other way on my computer